Describe the RM (Rate Monotonic) and EDF ( Early Deadline first) algorithm to schedule CPU with examples.
Answer #1

ANSWER:

Task model
a set of independent periodic tasks
EDF:
Whenever a new task arrive, sort the ready queue so that the task
closest to the end of its period assigned the highest priority
o Preempt the running task if it is not placed in the first of the queue
in the last sorting
FACT 1: EDF is optimal
EDF can schedule the task set if any one else can
FACT 2 (Scedulability test):
Σ Ci/Ti <= 1 iff the task set is schedulable

Example
Task set: {(2,5),(4,7)}
U = 2/5 + 4/7= 34/35 ~ 0.97 (schedulable!)

Example - Task set: {(2,5),(4,7)} FU = 2/5 + 4/7 = 34/35 v0.97 (schedulable!) | 西

EDF: + and –
Note that this is just the simple EDF algorithm; it works for all
types of tasks: periodic or non periodic
It is simple and works nicely in theory (+)
Simple schedulability test: U <= 1 (+)
Optimal (+)
Best CPU utilization (+)
Difficult to implement in practice. It is not very often adopted
due to the dynamic priority-assignment (expensive to sort the
ready queue on-line), which has nothing to do with the periods
of tasks. Note that Any task could get the highest priority (-)
Non stable: if any task instance fails to meet its deadline, the
system is not predictable, any instance of any task may fail (-)

Rate Monotonic Scheduling: task modelAssume a set of periodic tasks: (Ci,Ti)

Di=Ti

Tasks are always released at the start of their periods

Tasks are independent

RMS: fixed/static-priority scheduling

Rate Monotonic Fixed-Priority Assignment:o Tasks with smaller periods get higher priorities

Run-Time Scheduling:o Preemptive highest priority first

FACT: RMS is optimal in the sense:o If a task set is schedulable with any fixed-priorityscheduling algorithm, it is also schedulable with RMS

Example {(20,100),(40,150),(100,350)} Pr(T1)=1, Pr(T2)=2, Pr(T3)=3 T1 20 20 20 20 lo 100 200 300 40 40 150 40 300 o 40 30 350